NDOS Commissioning Meeting 2011/04/15¶
Attendance: Jon, Rick K., Alex, Ryan, Chris, Jan, Nick, Susan, Mat, Minerba, Tricia, Mark Messier, Gary, Denis, Enrique, Steve M., Alena, Ji Liu, Sue K., (and I've lost track of who else, sorry!)
- What Have We Learned During Commissioning?
- What Have We Accomplished During this Current Commissioning Period?
- What Still Needs to Be Done for NDOS?
- DataCheck Discussion
What Have We Learned During Commissioning?¶
- TECCs cause a great deal of noise when enabled
- APDs seem to be very sensitive to humidity (without cooling)
- When power is turned off for a long time and then turned back on, some APDs seem to produce a lot of noise.
- DSO scans
What Have We Accomplished During this Current Commissioning Period?¶
- We have seen NuMI and BNB neutrinos! Peaks in timing distributions and actual event displays.
- Successful data collections from 8 DCMs
- DCM hardware configuration read from and stored to central dB
- DAQ stability has improved greatly over the past few months:
- “3-hour DCM WOFO problem solved”
- DCM through-put increased
- Determination of individual pixel masks and thresholds
- Improved stability of underlying messaging service
- Resolved “lost” triggers problem
- Confirmed effectiveness of noise abatement boards on detector
What Still Needs to Be Done for NDOS?¶
- Run with cooled APDs while monitoring TECCs
- Integration between DAQ and DCS, especially with regard to DCMs
- Full DAQ configuration to the database
- Individual APD voltage settings
- Integrate pedestal run with normal DAQ run sequence
- Make TDUControl a RC Client
- DAQ partitioning
- Test coated APDs
- Automatic error recovery
- Jan presented one slide on DataCheck Present and Future
- Other suggestions:
- do not use BadChannel list
- record time of 1st event and last event (in subrun or run)
- record much of the summary information from DataCheck in central dB